On Saturday, Earl rode his bike for 2 hours longer than Ted. Earl traveled 135 km and Ted traveled 81km. If they both averaged the same rate of speed, how long did earl ride?
Answer by stanbon(75887)

On Saturday, Earl rode his bike for 2 hours longer than Ted. Earl traveled 135 km and Ted traveled 81km. If they both averaged the same rate of speed, how long did earl ride?
---------
Earl DATA:
distance = 135 km ; time = x+2 hrs ; rate = d/t = 135/(x+2) kph
-----------------
Ted DATA:
distance = 81 km ; time = x hrs ; rate = d/t = 81/x kph
----
Equation:
time = time
135/(x+2) = 81/x
----
135x = 81x + 162
---
54x = 162
x = 3 hrs (Ted riding time)
=====
Cheers,
Stan H.
------------